# Small-Signal Model for Inhomogeneous Helix Traveling-Wave Tubes using Transfer Matrices

### v1

@@ Abstract
We introduce a practical method for modeling the small-signal behavior of dispersive and inhomogeneous helix-type traveling-wave tube (TWT) amplifiers based on a generalization of the one-dimensional Pierce model. Our model is applicable to both single-stage and multi-stage TWTs. Like the Pierce model, we assume that electrons flow linearly in one direction, parallel and in proximity to a slow-wave structure (SWS) which guides a single dominant electromagnetic mode. Realistic helix TWTs are modeled with position-dependent and frequency-dependent SWS characteristics, such as loss, phase velocity, plasma frequency reduction factor, and interaction impedance. For the multi-stage helix TWT, we provide a simple lumped element circuit model for combining the stages separated by a sever, or gap, which attenuates the guided circuit mode while allowing the space-charge wave on the beam to pass freely to the next stage. The dispersive SWS characteristics are accounted for using full-wave eigenmode simulations for a realistic helix SWS supported by dielectric rods in a metal barrel, all of which contribute to the distributed circuit loss. We compare our computed gain vs frequency, computed using transfer matrices, to results found through particle-in-cell (PIC) simulations and the 1D TWT code LATTE to demonstrate the accuracy of our model. Furthermore, we demonstrate the ability of our model to reproduce gain ripple due to mismatches at the input and output ports of the TWT.
@@ I Introduction
For decades, traveling-wave tubes (TWTs) have been essential for applications which require high-power, broadband amplification at microwave and millimeter-wave frequencies. These applications include, but are not limited to radar, telecommunication, and electronic warfare. Traveling-wave tube amplifiers are highly reliable, long lasting, and efficient in their niche high frequency, high bandwidth, and high power applications, compared to their solid-state power amplifier counterparts (Chong and Menninger, 2010). The basic TWT amplifier is composed of a slow-wave structure (SWS) and a confined linear electron beam which flows in close proximity to the SWS. The guided modes that propagate in the SWS in the presence of the electron beam make the TWT a distributed amplifier by converting the kinetic energy of the electron beam into amplified waves when both the beam and the guided waves are velocity synchronized (Pierce, 1947, 1951). There are various types of SWS geometries for TWTs, though the one we consider here is the most common type: the tape helix. Helix-type TWTs can efficiently amplify waves over bandwidths of one octave and higher at moderate power levels, primarily limited by their capability to dissipate heat (Paoloni et al., 2021).
@@ II Problem Statement
The model described here is based on the generalization of the Pierce model in (Rouhi et al., 2021). However, unlike the works of (Pierce, 1947, 1950a, 1951; Rouhi et al., 2021) and the augmented model in (Rouhi et al., 2021), we make our model more general in several ways: by considering the dispersive coupling parameters as discussed in the next section, by modeling the EM properties the sever gap of the helix SWS using a capacitive network, and by considering real TWT features that are not necessarily homogeneous along $z$ (such as pitch tapering and distributed loss patterning).In Section III, we provide the analytical formulations for the system and transfer matrices, based on the generalized Pierce model, that can be used to model discretized segments of the “hot” TWT (i.e., with the electron beam present inside the SWS) as homogeneous equivalent transmission lines. A more detailed formulation can be found in (Rouhi et al., 2024). Using the unit cell dimensions provided in Section IV, we explain how the interaction impedance and attenuation coefficient used in our generalized Pierce model are computed using results from the eigenmode solver of the commercial software CST Studio Suite in Section V. In Section VI, we explain how we construct the equivalent transfer matrix for single-stage and two-stage helix TWTs. Then, in Section VII, we explain how we compute the small-signal gain and evolution of the state vector for the TWT examples, considering an equivalent lumped circuit network at the sever and terminations at the input/output ports of the TWT. Since our model relies on equivalent transmission lines (TLs) and load impedances, it allows us to observe how mismatches at the input and output ports of the TWT lead to gain ripple. In Section VIII, we show the gain vs frequency and evolution of the state vector versus position for the TWT examples and compare the results of our model to those from particle-in-cell (PIC) simulations and to those from the 1D TWT code LATTE. Finally, in Section IX, we explain how our model differs from the TWT code LATTE.
@@ III Generalized Pierce Model
To model TWTs that are inhomogeneous along their length, we segment each TWT stage of length $l$ into $S$ segments of homogeneous equivalent transmission lines (i.e. each segment has a constant attenuation, phase velocity, and characteristic impedance). The homogeneous equivalent transmission line model for a segment of the TWT stage is illustrated in Fig. 2. The homogeneous transmission line segments for each stage have a fixed length of $\Delta l=l/S$, where $S$ is the number of segments. Note that in this paper, $S$ does not need to be equal to the number of unit cells in the SWS, $N_{\mathrm{c}}$, which allows for more flexibility in the modeling of non homgenous parameters along the SWS. In the prior work by Rouhi et.al. (Rouhi et al., 2024), the TWT was partitioned into segments with length equal to the period of the serpentine waveguide unit cell (i.e., $\Delta l=d$ and $S=N_{\mathrm{c}}$). Specifically, subdividing the TWT stages by the geometric period of the SWS, $d$, works well for TWTs with no position-dependent features along their length. TWTs with pitch tapering and/or loss patterning sometimes need to be modeled with more transmission line segments than the number of unit cells (i.e. $\Delta l<d$) in order to adequately resolve these features (e.g. TWTs with have loss patterning near the sever that rapidly varies from one unit cell to the next). Table 1 summarizes the important parameters used to subdivide the TWT stages and define their physical length in our model (the subscripts 1 and 2 are later introduced to denote the parameters for the first and second stages in a two-stage TWT, respectively, in Sec. VI).
@@
Pierce theory is often used to describe the interaction between the space-charge waves of an electron beam and the guided EM mode using equivalent transmission lines in the small-signal regime with linearized equations. In Pierce theory, the TWT system is considered as a one-dimensional problem, with a guided EM wave on a SWS interacting with a linear space- and time-varying electron beam confined by a strong axial magnetic field, such that only electron movement in the axial ($+z$) direction is considered. For simplicity, we also assume that the axial electric field of the modes guided by the SWS and acting on the electron beam is approximately uniform over the cross-section of the electron beam, as was done in (Pierce, 1947, 1950a, 1951; Rouhi et al., 2021). The electron beam has an average equivalent kinetic voltage $V_{0}$ and average beam current $-I_{0}$ in the $+z$ direction (using conventional flow notation, positive values of $V_{0}$ and $I_{0}$ are used in the remainder of this paper). The average axial velocity of electrons in the beam is $u_{0}$ and it is approximately related to the average kinetic voltage as $V_{0}\approx u_{0}^{2}/\left(2\eta\right)$ for low-voltage beams (i.e. for $V_{0}$ near 10 kV or below (Jassem et al., 2020)), where $\eta=e/m_{0}=1.7588\times 10^{11}$ C/kg is the charge-to-rest-mass ratio of an electron. Note that this approximation for the electron kinetic voltage-velocity relation is needed, since the equations of the generalized Pierce model used here are derived using that approximation (Rouhi et al., 2021). The EM-charge modes that exist when the guided EM wave interacts with the electron beam are sometimes called “hot” modes, whereas passive EM modes guided by the SWS without the electron beam are called “cold” modes. For brevity, only the final equations necessary for the Pierce model are shown in this section. Detailed derivations of the equations governing the Pierce model may be found in the original Pierce papers (Pierce, 1947, 1950a, 1951), or in (Rouhi et al., 2021) and its supplementary material. Here, we augment the generalized Pierce model of (Rouhi et al., 2021, 2024) to include inhomogeneous features over the length of the TWT.
@@
The equivalent transmission line model for a homogeneous TWT segment is illustrated in Fig. 2(b). Under the small-signal, nonrelativistic approximations of the Pierce model, the four differential equations governing the dynamics of the electron beam and equivalent transmission line in the phasor domain can be represented in a compact matrix form at each frequency $\omega$ and longitudinal position $z$ along the circuit, with implicit $\exp(j\omega t)$ dependence, $\frac{d}{dz}\boldsymbol{\Psi}(z)=j\mathbf{\mathbf{\underline{M}_{\mathit{s}}}}\boldsymbol{\Psi}(z)$, with $z\in\left[(s-1)\Delta l,s\Delta l\right]$, where the subscript $s=1,...,S$ is the index corresponding to each segment along each stage of the TWT, the position-dependent $4\times 1$ state vector
@@
In Eqn. (2), $k_{\mathrm{c}}=\beta_{\mathrm{c}}-j\alpha_{\mathrm{c}}$ is the “cold” complex circuit wavenumber, that also accounts for losses in the SWS, and $Z_{\mathrm{c}}=V^{+}/I^{+}$ is the characteristic impedance of a given transmission line segment, $s$, in terms of the forward voltage and current waves traveling in the positive $z$-direction, denoted by the + sign (The total voltage and current in the presence of reflections is the sum of forward-traveling and backward-traveling waves, i.e. $V(z)=V^{+}(z)+V^{-}(z)$ and $I(z)=I^{+}(z)+I^{-}(z)$, see Ref. (Pozar, 2009, (Ch. 2))). Note that both the characteristic impedance and propagation constant in the matrix above can vary from segment to segment due to effects such as pitch tapering and position-dependent atteunation. For brevity, segment indices are not included in the subscripts for $k_{\mathrm{c}}$ and $Z_{\mathrm{c}}$. The definition of voltage and current on the helix SWS is chosen to yield the net EM power as $P=\frac{1}{2}\mathrm{Re}(VI^{*})$, where the asterisk (*) denotes complex conjugation. Here, we use the sheath helix model from (Kino and Paik, 1962; Paik, 1969) to define our helix voltage and current, and by neglecting losses in the sheath helix, we compute the resultant purely-real helix characteristic impedance $Z_{\mathrm{c}}$ in Appendix A (small losses are accounted for only by the complex propagation constant in our generalized Pierce model). Using a purely-real approximation for characteristic impedance and a complex propagation constant, is based on the simplifying assumption that the SWS is a low-loss transmission line (Collin, 1960, (Ch. 4))(Pozar, 2009, (Ch. 2))(Ramo, Whinnery, and Van Duzer, 1994, (Ch. 5)). The real part of the cold circuit wavenumber is related to the cold phase velocity of a segment (of length $\Delta l=l/S$, as explained above) of the low-loss SWS, $v_{\mathrm{c}}$, as $\beta_{\mathrm{c}}=\omega/v_{\mathrm{c}}$. Whereas the imaginary part of the propagation constant, $\alpha_{\mathrm{c}}$ is the attenuation per unit length (in Nepers per meter) for guided waves on the SWS. For each homogeneous segment of the equivalent transmission line that represents the SWS, the quantities $k_{\mathrm{c}}Z_{\mathrm{c}}$ and $k_{\mathrm{c}}/Z_{\mathrm{c}}$ in Eqn. (2) are related to the equivalent per-unit-length distributed series impedance and shunt admittance, as $Z=jk_{\mathrm{c}}Z_{\mathrm{c}}$ and $Y=jk_{\mathrm{c}}/Z_{\mathrm{c}}$ respectively, as illustrated in Fig. 2(b). Furthermore, the average electronic phase constant is $\beta_{0}=\omega/u_{0}$, $g=I_{0}\beta_{0}/(2V_{0})$, and $\zeta_{\mathrm{sc}}=2V_{0}\omega_{\mathrm{q}}^{2}/(\omega I_{0}u_{0})$, as was also considered in (Mealy, Abdelshafy, and Capolino, 2020; Rouhi et al., 2021, 2024). Note that $i_{\mathrm{s}}$ from Fig. 2 is represented in Eqn. (2) as $i_{\mathrm{s}}=-a(dI_{\mathrm{b}}/dz)=jagV_{\mathrm{b}}+ja\beta_{0}I_{\mathrm{b}}$. The frequency-dependent quantity $\omega_{\mathrm{q}}=R_{\mathrm{sc}}\omega_{\mathrm{p}}$ is the reduced plasma angular frequency of the electron beam, which is proportional to the plasma frequency $\omega_{\mathrm{p}}$ by the plasma frequency reduction factor $R_{\mathrm{sc}}$. The formulas used for $\omega_{\mathrm{p}}$ and $R_{\mathrm{sc}}$ are provided in Appendix E. The frequency-dependent coupling coefficient $a=(Z_{\mathrm{P}}/Z_{\mathrm{c}})^{1/2}$ is a factor which relates the actual characteristic impedance of the EM mode guided by the SWS to the beam-EM mode interaction impedance (sometimes referred to as Pierce impedance), as derived in (Rouhi et al., 2024). We also show, in Appendix B, that the system matrix $\mathbf{\underline{M}_{\mathit{s}}}$ may be transformed into $\mathbf{\underline{M}_{\mathit{s}}^{\prime}}$ to be in terms of interaction impedance $Z_{\mathrm{P}}$ only (see Appendix B), without dependence on the coupling coefficient $a$ or the characteristic impedance of the SWS $Z_{\mathrm{c}}$. In the original work by Pierce (Pierce, 1947, 1950a, 1950b, 1951), there was no distinction between these two impedances; $Z_{\mathrm{P}}$ was used as both the characteristic impedance and interaction impedance. However the equivalent voltages and currents on the transmission line for this transformed system matrix will also be scaled by $a$; one must consider the power of the guided mode when using the transformed model, rather than the helix voltage (and characteristic impedance) that can be defined in multiple ways (Henningsen, 1955). We note that, in this paper, the interaction impedance $Z_{\mathrm{P}}$ obtained by full-wave simulations is only used to determine the coupling coefficient $a$, unless one uses the alternative system matrix formulation provided in Appendix B. Generalized Pierce models using the coupling strength coefficient $a$ have also been developed in (Figotin and Reyes, 2013; Tamma and Capolino, 2014; Figotin, 2021a; Rouhi et al., 2021, 2024), but unlike those previous works, here we consider the fact that the TWT can be lossy, dispersive, and inhomogeneous along its length.

### v2

@@ Abstract
We introduce a practical method for modeling the small-signal behavior of frequency-dispersive and inhomogeneous helix-type traveling-wave tube (TWT) amplifiers based on a generalization of the one-dimensional Pierce model. Our model is applicable to both single-stage and multi-stage TWTs. Like the Pierce model, we assume that electrons flow linearly in one direction, parallel and in proximity to a slow-wave structure (SWS) which guides a single dominant electromagnetic mode. Realistic helix TWTs are modeled with position-dependent and frequency-dependent SWS characteristics, such as loss, phase velocity, plasma frequency reduction factor, interaction impedance, and the coupling factor that relates the SWS modal characteristic impedance to the interaction impedance. For the multi-stage helix TWT, we provide a simple lumped element circuit model for combining the stages separated by a sever, or gap, which attenuates the guided circuit mode while allowing the space-charge wave on the beam to pass freely to the next stage. The dispersive SWS characteristics are accounted for using full-wave eigenmode simulations for a realistic helix SWS supported by dielectric rods in a metal barrel, all of which contribute to the distributed circuit loss. We compare our computed gain vs frequency, computed using transfer matrices, to results found through particle-in-cell (PIC) simulations and the 1D TWT code LATTE to demonstrate the accuracy of our model. Furthermore, we demonstrate the ability of our model to reproduce gain ripple due to mismatches at the input and output ports of the TWT.
@@ I Introduction
For decades, traveling-wave tubes (TWTs) have been essential for applications which require high-power, broadband amplification at microwave and millimeter-wave frequencies. These applications include, but are not limited to radar, telecommunication, and electronic warfare. Traveling-wave tube amplifiers are highly reliable, long lasting, and efficient in their niche high frequency, high bandwidth, and high power applications, compared to their solid-state power amplifier counterparts (Chong and Menninger, 2010). The basic TWT amplifier is composed of a slow-wave structure (SWS) and a confined linear electron beam which flows in close proximity to the SWS. The guided modes that propagate in the SWS in the presence of the electron beam make the TWT a distributed amplifier by converting the kinetic energy of the electron beam into amplified waves when both the beam and the guided waves are velocity synchronized (Pierce, 1947, 1951). There are various types of SWS geometries for TWTs, though the one we consider here is the most common type: the tape helix. Helix-type TWTs can efficiently amplify waves over bandwidths of one octave and higher at moderate power levels, primarily limited by their capability to dissipate heat (Paoloni et al., 2021). Some notable examples of helix-type satellite TWTs with high efficiencies, long lifetimes, and moderate bandwidths can be found in Refs. (Bodmer et al., 1963; Chong and Menninger, 2010; Jacquez and Wada, 1986; Sauseng, Johnston, and Leigh, 1988; Srivastava et al., 1999; Wilson et al., 1991). Various wide bandwidth helix-type TWTs for applications such as electronic warfare have also been showcased in Refs. (Datta et al., 2009; Chong and Menninger, 2010; Frisoni, 1987; Braatz, Dohler, and Groshart, 1987).
@@ II Problem Statement
The model described here is based on the generalization of the Pierce model in (Rouhi et al., 2021). However, unlike the works of (Pierce, 1947, 1951, 1950) and the augmented model in (Rouhi et al., 2021), we make our model more general in several ways: by considering the dispersive coupling parameters as discussed in the next section, by modeling the EM properties the sever gap of the helix SWS using a capacitive network, and by considering real TWT features that are not necessarily homogeneous along $z$ (such as pitch tapering and distributed loss patterning).In Section III, we provide the analytical formulations for the system and transfer matrices, based on the generalized Pierce model, that can be used to model discretized segments of the “hot” TWT (i.e., with the electron beam present inside the SWS) as homogeneous equivalent transmission lines. A more detailed formulation can be found in (Rouhi et al., 2021, 2024). Furthermore, we highlight that the model shown in this work is compatible with TWTs with SWSs other than the helix geometry. For instance, ref. (Rouhi et al., 2024) demonstrates this linear model for a serpentine TWT, albeit without losses or position-dependent features. However, we restrict our discussion to the single-stage and two-stage helix-type TWT with arbitrary position-dependent loss distributions.
Using the unit cell dimensions provided in Section IV, we explain how the interaction impedance and attenuation coefficient used in our generalized Pierce model are computed using results from the eigenmode solver of the commercial software CST Studio Suite in Section V. In Section VI, we explain how we construct the equivalent transfer matrix for single-stage and two-stage helix TWTs. Then, in Section VII, we explain how we compute the small-signal gain and evolution of the state vector for the TWT examples, considering an equivalent lumped circuit network at the sever and terminations at the input/output ports of the TWT. Since our model relies on equivalent transmission lines (TLs) and load impedances, it allows us to observe how mismatches at the input and output ports of the TWT lead to gain ripple. In Section VIII, we show the gain vs frequency and evolution of the state vector versus position for the TWT examples and compare the results of our model to those from particle-in-cell (PIC) simulations and to those from the 1D TWT code LATTE. Finally, in Section IX, we explain how our model differs from the TWT code LATTE.
@@ III Generalized Pierce Model
To model TWTs that are inhomogeneous along their length, we segment each TWT stage of length $l$ into $S$ segments of homogeneous equivalent transmission lines (i.e. each segment has a constant attenuation, phase velocity, and characteristic impedance), similarly to what has been done in (Chernin et al., 2012). The homogeneous equivalent transmission line model for a segment of the TWT stage is illustrated in Fig. 2. The homogeneous transmission line segments for each stage have a fixed length of $\Delta l=l/S$, where $S$ is the number of segments. Note that in this paper, $S$ does not need to be equal to the number of unit cells in the SWS, $N_{\mathrm{c}}$, which allows for more flexibility in the modeling of non homgenous parameters along the SWS. In the prior work by Rouhi et.al. (Rouhi et al., 2024), the TWT was partitioned into segments with length equal to the period of the serpentine waveguide unit cell (i.e., $\Delta l=d$ and $S=N_{\mathrm{c}}$). Specifically, subdividing the TWT stages by the geometric period of the SWS, $d$, works well for TWTs with no position-dependent features along their length. TWTs with pitch tapering and/or loss patterning sometimes need to be modeled with more transmission line segments than the number of unit cells (i.e. $\Delta l<d$) in order to adequately resolve these features (e.g. TWTs with have loss patterning near the sever that rapidly varies from one unit cell to the next). Table 1 summarizes the important parameters used to subdivide the TWT stages and define their physical length in our model (the subscripts 1 and 2 are later introduced to denote the parameters for the first and second stages in a two-stage TWT, respectively, in Sec. VI).
@@
Pierce theory is often used to describe the interaction between the space-charge waves of an electron beam and the guided EM mode using equivalent transmission lines in the small-signal regime with linearized equations. In Pierce theory, the TWT system is considered as a one-dimensional problem, with a guided EM wave on a SWS interacting with a linear space- and time-varying electron beam confined by a strong axial magnetic field, such that only electron movement in the axial ($+z$) direction is considered. For simplicity, we also assume that the axial electric field of the modes guided by the SWS and acting on the electron beam is approximately uniform over the cross-section of the electron beam, as was done in (Pierce, 1947, 1951; Rouhi et al., 2021; Pierce, 1950). The electron beam has an average equivalent kinetic voltage $V_{0}$ and average beam current $-I_{0}$ in the $+z$ direction (using conventional flow notation, positive values of $V_{0}$ and $I_{0}$ are used in the remainder of this paper). The average axial velocity of electrons in the beam is $u_{0}$ and it is approximately related to the average kinetic voltage as $V_{0}\approx u_{0}^{2}/\left(2\eta\right)$ for low-voltage beams (i.e. for $V_{0}$ near 10 kV or below (Jassem et al., 2020)), where $\eta=e/m_{0}=1.7588\times 10^{11}$ C/kg is the charge-to-rest-mass ratio of an electron. Note that this approximation for the electron kinetic voltage-velocity relation is needed, since the equations of the generalized Pierce model used here are derived using that approximation (Rouhi et al., 2021). For higher beam voltages, it is necessary to account for changes in the effective electron mass in the derivation of the electron beam equations of (Rouhi et al., 2021). Development of a generalized Pierce model with relativistic corrections is planned for a future paper. Hence, we limit our generalized Pierce model in this work to the non-relativistic regime, as it builds upon our previous works (which were also formulated using non-relativistic equations) (Figotin and Reyes, 2013; Tamma and Capolino, 2014; Figotin, 2021a; Rouhi et al., 2021, 2024). The EM-charge modes that exist when the guided EM wave interacts with the electron beam are sometimes called “hot” modes, whereas passive EM modes guided by the SWS without the electron beam are called “cold” modes. For brevity, only the final equations necessary for the Pierce model are shown in this section. Detailed derivations of the equations governing the Pierce model may be found in the original Pierce papers (Pierce, 1947, 1951, 1950), or in (Rouhi et al., 2021) and its supplementary material. Here, we augment the generalized Pierce model of (Rouhi et al., 2021, 2024) to include inhomogeneous features over the length of the TWT.
@@
The equivalent transmission line model for a homogeneous TWT segment is illustrated in Fig. 2(b). Under the small-signal, nonrelativistic approximations of the Pierce model, the four differential equations governing the dynamics of the electron beam and equivalent transmission line in the phasor domain can be represented in a compact matrix form at each frequency $\omega$ and longitudinal position $z$ along the circuit, with implicit $\exp(j\omega t)$ dependence, $\frac{d}{dz}\boldsymbol{\Psi}(z)=-j\mathbf{\mathbf{\underline{M}_{\mathit{s}}}}\boldsymbol{\Psi}(z)$, with $z\in\left[(s-1)\Delta l,s\Delta l\right]$, where the subscript $s=1,...,S$ is the index corresponding to each segment along each stage of the TWT, the position-dependent $4\times 1$ state vector
@@
In Eqn. (2), $k_{\mathrm{c}}=\beta_{\mathrm{c}}-j\alpha_{\mathrm{c}}$ is the “cold” complex circuit wavenumber, that also accounts for losses in the SWS, and $Z_{\mathrm{c}}=V^{+}/I^{+}$ is the characteristic impedance of a given transmission line segment, $s$, in terms of the forward voltage and current waves traveling in the positive $z$-direction, denoted by the + sign (The total voltage and current in the presence of reflections is the sum of forward-traveling and backward-traveling waves, i.e. $V(z)=V^{+}(z)+V^{-}(z)$ and $I(z)=I^{+}(z)+I^{-}(z)$, see Ref. (Pozar, 2009, (Ch. 2))). Note that both the characteristic impedance and propagation constant in the matrix above can vary from segment to segment due to effects such as pitch tapering and position-dependent atteunation. For brevity, segment indices are not included in the subscripts for $k_{\mathrm{c}}$ and $Z_{\mathrm{c}}$. The definition of voltage and current on the helix SWS is chosen to yield the net EM power as $P=\frac{1}{2}\mathrm{Re}(VI^{*})$, where the asterisk (*) denotes complex conjugation. Here, we use the sheath helix model from (Kino and Paik, 1962; Paik, 1969) to define our helix voltage and current, and by neglecting losses in the sheath helix, we compute the resultant purely-real helix characteristic impedance $Z_{\mathrm{c}}$ in Appendix A (small losses are accounted for only by the complex propagation constant in our generalized Pierce model). Using a purely-real approximation for characteristic impedance and a complex propagation constant, is based on the simplifying assumption that the SWS is a low-loss transmission line (Collin, 1960, (Ch. 4))(Pozar, 2009, (Ch. 2))(Ramo, Whinnery, and Van Duzer, 1994, (Ch. 5)). The real part of the cold circuit wavenumber is related to the cold phase velocity of a segment (of length $\Delta l=l/S$, as explained above) of the low-loss SWS, $v_{\mathrm{c}}$, as $\beta_{\mathrm{c}}=\omega/v_{\mathrm{c}}$. Whereas the imaginary part of the propagation constant, $\alpha_{\mathrm{c}}$ is the attenuation per unit length (in Nepers per meter) for guided waves on the SWS. For each homogeneous segment of the equivalent transmission line that represents the SWS, the quantities $k_{\mathrm{c}}Z_{\mathrm{c}}$ and $k_{\mathrm{c}}/Z_{\mathrm{c}}$ in Eqn. (2) are related to the equivalent per-unit-length distributed series impedance and shunt admittance, as $Z=jk_{\mathrm{c}}Z_{\mathrm{c}}$ and $Y=jk_{\mathrm{c}}/Z_{\mathrm{c}}$ respectively, as illustrated in Fig. 2(b). Furthermore, the average electronic phase constant is $\beta_{0}=\omega/u_{0}$, $g=I_{0}\beta_{0}/(2V_{0})$, and $\zeta_{\mathrm{sc}}=2V_{0}\omega_{\mathrm{q}}^{2}/(\omega I_{0}u_{0})$, as was also considered in (Mealy, Abdelshafy, and Capolino, 2020; Rouhi et al., 2021, 2024). Note that $i_{\mathrm{s}}$ from Fig. 2 is represented in Eqn. (2) as $i_{\mathrm{s}}=-a(dI_{\mathrm{b}}/dz)=jagV_{\mathrm{b}}+ja\beta_{0}I_{\mathrm{b}}$. The frequency-dependent quantity $\omega_{\mathrm{q}}=R_{\mathrm{sc}}\omega_{\mathrm{p}}$ is the reduced plasma angular frequency of the electron beam, which is proportional to the plasma frequency $\omega_{\mathrm{p}}$ by the plasma frequency reduction factor $R_{\mathrm{sc}}$. The formulas used for $\omega_{\mathrm{p}}$ and $R_{\mathrm{sc}}$ are provided in Appendix E. The frequency-dependent coupling coefficient $a=(Z_{\mathrm{P}}/Z_{\mathrm{c}})^{1/2}$ is a factor which relates the actual characteristic impedance of the EM mode guided by the SWS to the beam-EM mode interaction impedance (sometimes referred to as Pierce impedance), as derived in (Rouhi et al., 2024). We also show, in Appendix B, that the system matrix $\mathbf{\underline{M}_{\mathit{s}}}$ may be transformed into $\mathbf{\underline{M}_{\mathit{s}}^{\prime}}$ to be in terms of interaction impedance $Z_{\mathrm{P}}$ only (see Appendix B), without dependence on the coupling coefficient $a$ or the characteristic impedance of the SWS $Z_{\mathrm{c}}$. In the original work by Pierce (Pierce, 1947, 1951, 1950), there was no distinction between these two impedances; $Z_{\mathrm{P}}$ was used as both the characteristic impedance and interaction impedance. However the equivalent voltages and currents on the transmission line for this transformed system matrix will also be scaled by $a$; one must consider the power of the guided mode when using the transformed model, rather than the helix voltage (and characteristic impedance) that can be defined in multiple ways (Henningsen, 1955). We note that, in this paper, the interaction impedance $Z_{\mathrm{P}}$ obtained by full-wave simulations is only used to determine the coupling coefficient $a$, unless one uses the alternative system matrix formulation provided in Appendix B. Generalized Pierce models using the coupling strength coefficient $a$ have also been developed in (Figotin and Reyes, 2013; Tamma and Capolino, 2014; Figotin, 2021a; Rouhi et al., 2021, 2024), but unlike those previous works, here we consider the fact that the TWT can be lossy, dispersive, and inhomogeneous along its length. Furthermore, we demonstrate in Appendix F that the fourth-order modal dispersion relation from Pierce theory (Pierce, 1951, 1947)(Pierce, 1950, Ch. 2)(Tsimring, 2006, Ch. 8) can be obtained using the system matrix of Eqn. (2), provided that the characteristic impedance $Z_{\mathrm{c}}$is assumed equal to the interaction impedance $Z_{\mathrm{P}}$(as is done in conventional Pierce theory) to make the coupling coefficient $a$ equal to unity. We also note that, unlike with our model, the original works by Pierce did not consider dispersive parameters (e.g. frequency-dependent interaction impedance and phase velocity for the modes supported by the cold SWS) in solving for the modal dispersion relation for the TWT (Pierce, 1951, 1950, 1947).